# FlowVLA: Thinking in Motion with a Visual Chain of Thought

### v2

@@ 3.2 Main Results on Robotics Benchmarks (Q1)
Table 1: Results on the LIBERO Benchmark. We report the final task success rate (%). We compare FlowVLA against state-of-the-art methods, grouped by their core methodology. Our key comparison is within the w/ World Model group, where our Visual CoT pre-training demonstrates superior performance. Notably, FlowVLA achieves this without relying on massive external datasets, highlighting the efficiency of our proposed framework.
@@
Diffusion Policy Chi et al. (2023) $\times$ 78.3 92.5 68.3 50.5 72.4
Octo Team et al. (2024) $\checkmark$ 78.9 85.7 84.6 51.1 75.1
OpenVLA Kim et al. (2024) $\checkmark$ 84.7 88.4 79.2 53.7 76.5
DiT Policy Hou et al. (2025) $\checkmark$ 84.2 96.3 85.4 63.8 82.4
TraceVLA Zheng et al. (2024) $\checkmark$ 84.6 85.2 75.1 54.1 74.8
SpatialVLA Qu et al. (2025) $\checkmark$ 88.2 89.9 78.6 55.5 78.1
pi0-FAST Pertsch et al. (2025) $\checkmark$ 96.4 96.8 88.6 60.2 85.5
ThinkAct Huang et al. (2025a) $\checkmark$ 88.3 91.4 87.1 70.9 84.4
@@
WorldVLA Cen et al. (2025) $\times$ 85.6 89.0 82.6 59.0 79.1
UniVLA Wang et al. (2025) $\times$ 92.6 93.8 86.6 63.0 84.0
CoT-VLA Zhao et al. (2025) $\checkmark$ 87.5 91.6 87.6 69.0 81.1
FlowVLA (ours) $\times$ 93.2 95.0 91.6 72.6 88.1
† Our reported UniVLA result is from our re-implementation, pre-trained only on LIBERO without wrist camera images for a fair comparison. The original paper utilized large-scale robotics data and an additional wrist camera.
Table 2: Results on the SimplerEnv-WidowX benchmark. We report the final task success rate (%). FlowVLA significantly surpasses prior methods, demonstrating superior robustness to the visual domain shifts present in this benchmark. Best results are in bold.
@@ 3.3 Analysis of World Modeling Capabilities (Q2)
To demonstrate the superiority of our Visual Chain of Thought (Visual CoT) framework, we conduct a detailed qualitative analysis on the challenging, real-world Bridge V2 dataset. Our investigation reveals two distinct and critical failure modes in the standard next-frame prediction baseline.
First, Figure 3 highlights failures in physical plausibility. In these examples, the baseline model generates physically incoherent rollouts, such as causing the robotic arm to suddenly vanish or producing inconsistent object motion. This indicates a fundamental inability to model the basic physical continuity of a scene. Second, Figure 4 illustrates a more subtle but equally critical issue: semantic inconsistency. Here, while the predicted frames from the baseline may appear visually coherent, the depicted action fails to follow the given language command. This reveals a disconnect between language understanding and visual forecasting.
In stark contrast, FlowVLA successfully navigates both challenges across all scenarios. By first reasoning about motion dynamics via optical flow, our model generates predictions that are not only physically plausible but also semantically aligned with the task instructions, showcasing the robustness and generalizability of our approach.
@@
Figure 3: Analysis of Physical Plausibility on the Bridge V2 Dataset. This figure highlights common physical failures in the baseline model. In both examples, the baseline model (middle row) struggles to maintain physical consistency, leading to implausible outcomes such as a disappearing manipulator or erratic object behavior. In contrast, FlowVLA (bottom row), guided by its motion-first reasoning, produces stable and physically coherent predictions that accurately reflect the scene’s dynamics.
@@
Figure 4: Analysis of Semantic Alignment on the Bridge V2 Dataset. This figure illustrates the baseline’s failure to align predictions with language instructions. While the predicted frames from baseline model (middle row) might appear visually plausible at a glance, the resulting motion does not correspond to the specified task (e.g., moving an object in the wrong direction). FlowVLA (bottom row) again demonstrates superior performance, correctly interpreting the command and generating a corresponding visual trajectory. This underscores that our Visual CoT not only improves physical realism but also enhances the model’s ability to ground language in action.
@@ 3.4 Convergence Speed and Data Efficiency(Q3)
Figure 5 illustrates FlowVLA’s dramatic advantage in training and sample efficiency. In the full-data setting (Figure 5a), FlowVLA proves vastly more sample-efficient, reaching the baseline’s peak performance (0.64) with only one-third of the training steps (2k vs. 6k) while also achieving a higher final success rate of 0.73.
This efficiency advantage is amplified in the more challenging low-data regime (Figure 5b). Here, the performance gap widens substantially. FlowVLA not only achieves a 55% higher peak success rate relative to the baseline (0.48 vs. 0.31) but also surpasses the baseline’s peak performance in just 1,000 steps. This substantial improvement in sample efficiency validates our core hypothesis: by requiring the model to explicitly reason about motion via a visual chain-of-thought, FlowVLA benefits from a powerful inductive bias. This simplifies the learning of physical dynamics from raw pixels, leading to a more effective and robust learning process, particularly when data is limited.
@@ 3.5 Ablation Studies (Q4)
We first remove the entire Visual Chain-of-Thought (CoT) structure, which causes our model to degenerate into the UniVLA baseline. As shown in Table 3, the success rate drops sharply from 73.0% to 64.0%. This significant 9-point drop confirms that the explicit, step-by-step reasoning process, where the model first thinks about "how to move" before predicting the outcome, is the primary driver of our model’s performance gain.
@@
Finally, we challenge the core design of interleaving information. We restructure the input sequence by grouping all visual frames first, followed by all corresponding flow frames (i.e., $v_{0},v_{1},...,f_{0},f_{1},$...). This configuration leads to a severe performance collapse, with the success rate plummeting to 49.4%. This is because the model can no longer leverage the predicted motion $f_{t}$ to inform the generation of the next state $v_{t+1}$ in a causal, forward-looking manner. This result provides strong evidence that the “interleaved, step-by-step causal chain” ($v_{t}\rightarrow f_{t}\rightarrow v_{t+1}$) is essential for effective planning and action generation.
@@ 4.1 Vision-Language-Action (VLA) Models
The dominant paradigm for creating generalist robot agents is the Vision-Language-Action (VLA) model Zitkovich et al. (2023), Kim et al. (2024), Black et al. (2024), Song et al. (2025b, c). These models extend large, pre-trained Vision-Language Models (VLMs) by fine-tuning them on extensive robotics datasets O’Neill et al. (2024). Architectures like RT-2 Zitkovich et al. (2023), CEED-VLA Song et al. (2025a) and OpenVLA Kim et al. (2024) treat action generation as a sequence modeling problem, directly mapping visual and textual inputs to discretized action tokens. Other recent works have focused on improving the action representation itself, using techniques like diffusion policies Chi et al. (2023) or flow matching Black et al. (2024). While this end-to-end approach has demonstrated remarkable generalization, it often treats the environment’s physical dynamics as a “black box”. The policy is learned reactively, without an explicit, underlying model of how the world functions or evolves. FlowVLA diverges from this standard VLA formulation by prioritizing world understanding over immediate action generation. Its pre-training objective is not to learn a policy ($v_{t}\rightarrow a_{t}$), but to build a robust world model by learning the physical transition function of the environment ($v_{t}\rightarrow v_{t+1}$). This “dynamics-first” approach establishes a solid foundation of physical knowledge before it is adapted for downstream control.
@@ 4.2 World Models for Robotics
The concept of a world model, which learns a model of its environment to plan or imagine future outcomes Ha and Schmidhuber (2018), is increasingly vital in robotics. Recent works have leveraged this idea for policy learning. For example, some models use video prediction as a form of self-supervised pre-training to improve downstream task performance Wang et al. (2025), Wu et al. (2023). Others, like WorldVLA Cen et al. (2025), propose architectures that jointly learn to predict both the next frame and the next action, creating a tight loop between prediction and control. A common thread in these approaches is the direct prediction of the next frame, modeling the transition as $v_{t}\rightarrow v_{t+1}$. However, this direct objective forces a single network to simultaneously handle two distinct problems: understanding static scene properties (appearance, texture, lighting) and modeling complex physical dynamics (motion, interaction, causality). This entanglement can result in inefficient learning and physically implausible predictions, such as blurry or distorted futures. In contrast, FlowVLA avoids this entanglement with its Visual Chain of Thought framework. We decompose the prediction into a “frame $\rightarrow$ flow $\rightarrow$ frame” reasoning process. By forcing the model to first predict an intermediate optical flow field ($f_{t}$), we explicitly decouple the learning of dynamics (how things move) from appearance (what they look like), resulting in a more causally-grounded world model.
@@ 5 Conclusion
We proposed the Visual Chain of Thought (Visual CoT) as a new principle for world model learning, instantiated in FlowVLA. By decomposing prediction into a motion–then–appearance process, FlowVLA learns more coherent dynamics within a unified autoregressive model. Experiments on robotics benchmarks demonstrate state-of-the-art performance and improved sample efficiency, highlighting the value of explicit motion reasoning for bridging perception and control.

### v3

@@ 3.2 Evaluations Results (Q1)
Results on Real-Robot. As shown in Table 3, FlowVLA exhibits clear advantages over all baselines across the four real-world Cobot tasks, including both single-arm and bimanual operations. The improvements are particularly pronounced in more complex bimanual tasks such as Placing two cola cans into a box and Lifting a pot using both arms, where precise coordination and dynamic interaction are required. During evaluation, each task was executed 25 times to ensure statistically reliable success rates. These results confirm that FlowVLA’s explicit motion reasoning and world model-based long-horizon planning are effective in real-world settings, enabling rapid adaptation to the specific embodiment and visual conditions of the Cobot platform.
Table 1: Results on the LIBERO Benchmark Liu et al. (2023). We report the final task success rate (%) and compare FlowVLA against state-of-the-art methods, grouped by their core methodology. The results demonstrate that our Visual CoT pre-training leads to superior performance, highlighting the efficiency of our proposed framework.
@@
Diffusion Policy Chi et al. (2023) 78.3 92.5 68.3 50.5 72.4
Octo Team et al. (2024) 78.9 85.7 84.6 51.1 75.1
OpenVLA Kim et al. (2024b) 84.7 88.4 79.2 53.7 76.5
DiT Policy Hou et al. (2025) 84.2 96.3 85.4 63.8 82.4
TraceVLA Zheng et al. (2024) 84.6 85.2 75.1 54.1 74.8
SpatialVLA Qu et al. (2025) 88.2 89.9 78.6 55.5 78.1
pi0-FAST Pertsch et al. (2025) 96.4 96.8 88.6 60.2 85.5
ThinkAct Huang et al. (2025a) 88.3 91.4 87.1 70.9 84.4
@@
WorldVLA Cen et al. (2025) 85.6 89.0 82.6 59.0 79.1
[…]
CoT-VLA Zhao et al. (2025) 87.5 91.6 87.6 69.0 81.1
FlowVLA (ours) 93.2 95.0 91.6 72.6 88.1
† Our reported UniVLA result is from our re-implementation, pre-trained only on LIBERO without wrist camera images for a fair comparison.
Table 2: Results on the SimplerEnv-WidowX benchmark Li et al. (2024). We report the final task success rate (%). FlowVLA significantly surpasses prior methods, demonstrating superior robustness to the visual domain shifts present in this benchmark. Best results are in bold.
@@
(a) Platform setup
(b) Example tasks
Figure 3: AgileX Cobot dual-arm platform and real-world manipulation tasks: (a) system setup: leader arms are user-operated, follower arms mirror the actions. Vision is provided by a front camera for global scene view and wrist cameras for close-up workspace observation. (b) representative single-arm and bimanual operations: from simple single-arm tasks to complex long-horizon bimanual manipulations.
Table 3: Results on the Real-world AgileX Cobot platform. We report the task success rates (%). The evaluation covers four manipulation tasks of varying difficulty, from single-arm to bimanual operation. Best results are in bold.
Model Stack Bowls Place Vegetable Place Bottles Lift Pot Avg.
ACT Zhao et al. (2023) 32.0 24.0 12.0 8.0 19.0
OpenVLA Kim et al. (2024b) 28.0 20.0 20.0 12.0 20.0
UniVLA Wang et al. (2025) 48.0 40.0 16.0 20.0 31.0
FlowVLA (Ours) 56.0 60.0 32.0 28.0 44.0
@@ 3.3 Analysis of World Modeling Capabilities (Q2)
To demonstrate the superiority of our motion reasoning framework compared with approaches that learn world dynamics implicitly, we conduct a detailed qualitative analysis on the challenging, real-world Bridge V2 dataset. The standard next-frame prediction baseline suffers from two distinct and critical failure modes.
  • Failures in Physical Plausibility. As highlighted in Figure 4, the baseline model generates physically incoherent rollouts, such as causing the robotic arm to suddenly vanish or producing inconsistent object motion. This indicates a fundamental inability to model the basic physical continuity of a scene.
  • Semantic Inconsistency. Figure 5 illustrates a more subtle but equally critical issue. While the predicted frames from the baseline may appear visually coherent, the depicted action fails to follow the given language command. This reveals a disconnection between language understanding and visual forecasting.
In contrast, FlowVLA successfully overcomes the above challenges. By first reasoning about motion dynamics via optical flow, our model generates predictions that are not only physically plausible but also semantically aligned with the task instructions, showcasing the robustness and generalizability of our approach.
@@
Figure 4: Analysis of Physical Plausibility on the Bridge V2 Dataset. This figure highlights common physical failures in the baseline model. In both examples, the baseline model (top row) struggles to maintain physical consistency, leading to implausible outcomes such as a disappearing manipulator or erratic object behavior. In contrast, FlowVLA (bottom row), guided by its motion-first reasoning, produces stable and physically coherent predictions that accurately reflect the scene’s dynamics.
@@
Figure 5: Analysis of Semantic Alignment on the Bridge V2 Dataset. This figure illustrates the baseline’s failure to align predictions with language instructions. While the predicted frames from baseline model (top row) might appear visually plausible at a glance, the resulting motion does not correspond to the specified task (e.g., moving an object in the wrong direction). FlowVLA (bottom row) again demonstrates superior performance, correctly interpreting the command and generating a corresponding visual trajectory. This underscores that our Visual CoT not only improves physical realism but also enhances the model’s ability to ground language in action.
@@ 3.4 Convergence Speed and Data Efficiency(Q3)
To isolate and evaluate the impact of our Visual CoT pre-training, we conduct a direct comparison between FlowVLA and its foundational baseline, UniVLA. Figure 6 illustrates FlowVLA’s dramatic advantage in training and sample efficiency. In the full-data setting (Figure 6(a)), FlowVLA proves vastly more sample-efficient, reaching the baseline’s peak performance (0.64) with only one-third of the training steps (2k vs. 6k) while also achieving a higher final success rate of 0.73.
This efficiency advantage is amplified in the more challenging low-data regime (Figure 6(b)). Here, the performance gap widens substantially. FlowVLA not only achieves a 55% higher peak success rate relative to the baseline (0.48 vs. 0.31) but also surpasses the baseline’s peak performance in just 1k steps. This substantial improvement in sample efficiency validates our core hypothesis: by requiring the model to explicitly reason about motion via a visual chain-of-thought, FlowVLA benefits from a powerful inductive bias. This simplifies the learning of physical dynamics from raw pixels, leading to a more effective and robust learning process, particularly when data is limited.
@@ 3.5 Ablation Studies (Q4)
We first remove the entire Visual Chain-of-Thought (CoT) structure, which causes our model to degenerate into the UniVLA baseline. As shown in Table 4, the success rate drops sharply from 73.0% to 64.0%. This significant 9-point drop confirms that the explicit, step-by-step reasoning process, where the model first thinks about “how to move” before predicting the outcome, is the primary driver of our model’s performance gain.
@@
Finally, we validate the core design of interleaving information. We restructure the input sequence into the format $v_{0},v_{1},\dots,f_{0},f_{1},\dots$, where all visual frames are grouped first, followed by all corresponding flow frames. This configuration leads to a severe performance collapse, with the success rate plummeting to 49.4%. This is because the model can no longer leverage the predicted motion $f_{t}$ to inform the generation of the next state $v_{t+1}$ in a causal, forward-looking manner. This result provides strong evidence that the “interleaved, step-by-step causal chain” ($v_{t}\rightarrow f_{t}\rightarrow v_{t+1}$) is essential for effective planning and action generation.
@@ 4.1 Vision-Language-Action (VLA) Models
The dominant paradigm for creating generalist robot agents is the Vision-Language-Action (VLA) model Zitkovich et al. (2023), Kim et al. (2024b), Black et al. (2024), Song et al. (2025b, c). These models extend large, pre-trained Vision-Language Models (VLMs) by fine-tuning them on extensive robotics datasets O’Neill et al. (2024). Architectures like RT-2 Zitkovich et al. (2023), CEED-VLA Song et al. (2025a) and OpenVLA Kim et al. (2024b) treat action generation as a sequence modeling problem, directly mapping visual and textual inputs to discretized action tokens. Other recent works have focused on improving the action representation itself, using techniques like diffusion policies Chi et al. (2023) or flow matching Black et al. (2024). While this end-to-end approach has demonstrated remarkable generalization, it often treats the environment’s physical dynamics as a “black box”. The policy is learned reactively, without an explicit, underlying model of how the world functions or evolves. FlowVLA diverges from this standard VLA formulation by prioritizing world understanding over immediate action generation. Its pre-training objective is not to learn a policy ($v_{t}\rightarrow a_{t}$), but to build a robust world model by learning the physical transition function of the environment ($v_{t}\rightarrow v_{t+1}$). This “dynamics-first” approach establishes a solid foundation of physical knowledge before it is adapted for downstream control.
@@ 4.2 World Models for Robotics
The concept of a world model, which learns a model of its environment to plan or imagine future outcomes Ha and Schmidhuber (2018), is increasingly vital in robotics. Recent works have leveraged this idea for policy learning. For example, some models use video prediction as a form of self-supervised pre-training to improve downstream task performance Wang et al. (2025), Wu et al. (2023). Others, like WorldVLA Cen et al. (2025), propose architectures that jointly learn to predict both the next frame and the next action, creating a tight loop between prediction and control. A common thread in these approaches is the direct prediction of the next frame, modeling the transition as $v_{t}\rightarrow v_{t+1}$. However, this direct objective forces a single network to simultaneously handle two distinct problems: understanding static scene properties (appearance, texture, lighting) and modeling complex physical dynamics (motion, interaction, causality). This entanglement can result in inefficient learning and physically implausible predictions, such as blurry or distorted futures. In contrast, FlowVLA avoids this entanglement with its Visual Chain of Thought framework. We decompose the prediction into a “frame flow frame” reasoning process. By forcing the model to first predict an intermediate optical flow field ($f_{t}$), we explicitly decouple the learning of dynamics (how things move) from appearance (what they look like), resulting in a more causally-grounded world model.
@@ 5 Conclusion
We proposed the Visual Chain of Thought (Visual CoT) as a new paradigm for world model learning, instantiated in FlowVLA. By decomposing prediction into an explicit motion–then–appearance reasoning sequence $v_{t}\rightarrow f_{t}\rightarrow v_{t+1}$, our model learns physically grounded representations that align with the demands of downstream action prediction. We introduce a two-stage training paradigm. At the first stage, we pre-train the world model with Visual CoT to build motion-coherent and physically plausible dynamics knowledge from video data. At the second stage, we fine-tune the policy to adapt this knowledge to generate precise robot actions. This design directly addresses the gap between pre-training and fine-tuning in VLA models, leading to improved sample efficiency and robust performance. Experiments on simulation and real-world manipulation benchmarks validate the effectiveness of this motion-first approach, confirming that explicit motion reasoning is a powerful inductive bias for bridging perception and control in generalist robotics.